A rectangular pool has dimensions of 40 ft. and 60 ft. If the pool has a patio around it that equals the area, how wide is the distance from the pool edge to the patio edge?
I got the answer x=10, however the question ask the distance from the pool edge to the patio edge?
What is the answer? 80?

Answer by mananth(16949)

A rectangular pool has dimensions of 40 ft. and 60 ft. If the pool has a patio around it that equals the area, how wide is the distance from the pool edge to the patio edge
Area of pool = 40*60 = 2400 sq.ft
let the width of the patio be x
width of pool and patio = 40+2x
length = 60+2x
Total area - area of pool = 2400
(40+2x)(60+2x)-2400=2400
2400+80x+120x+4x^2-2400=2400
4x^2+200x-2400=0
/4
x^2+50x-600=0
x^2+60x-10x-600=0
x(x+60)-10(x+60)=0
(x-10)(x+60)=0
x=10 ft the width
you are correct